OpenStudy (anonymous):

What is the simplest form of:

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{x ^{3}} }{ \sqrt[5]{x ^{2}} }\]

OpenStudy (gorv):

@AngelCriner in numerator is it3 or 8

OpenStudy (anonymous):

It's 3

OpenStudy (gorv):

(x^2)^(1/3)

OpenStudy (gorv):

\[\frac{ x ^{\frac{ 3 }{ 8 }}}{x^{\frac{ 3 }{ 5 }}}\]

OpenStudy (anonymous):

Do I have to find a common denominator?

OpenStudy (gorv):

both have common base x

OpenStudy (anonymous):

or are they the same... looks like an 8... or a 5

OpenStudy (anonymous):

What does that mean?

OpenStudy (gorv):

just find out power of x in numerator and denominator

OpenStudy (gorv):

actually power arenot clearly readable in your Q

OpenStudy (anonymous):

I just typed it from the question... it's definitely a 3... the text is just so tiny on here. I honestly don't know how to do that... I'm not very good at math or fluen in it's vocabulary.